Can someone please help me with this problem:
Find the compound amount for the deposit. Round to the nearest cent.
$5000 at 7% compounded annually for 6 years.

Answer by rfer(16322)

FV=P(1+r)^n
FV=5000(1+.07)^6
FV=5000(1.50073)
FV=$7503.65